Find the square of the following numbers. 
46
Find the square of the following numbers. 
46

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

(46)^2
=(50-4)^2
= (50)^2+ (4)^2- 2×50×4 [Since, (a+b)^2= a^2+b^2+2ab]
=2500+16–400
= 2116

​Thanks